l vodafone AU 3:21 pm KAssignment 1 Problems Fillable.pdf.. Prohlom 2 3 of 5 uns law for an economy is given by the following equation 100 (부)--2(u_u*) 2.1 Briefly explain the meaning of Okuns law 2.2 Use Okuns law to complete the missing data in the following table. Natural unemployment rate 6 PotentiaUnemployment ear Real GDP GDP rate 2015 7840 8000 8100 2017 8200 4.5 2017 8415 8250
Answer #1

Okun's law suggests a relationship between output gap and unemployment gap in which for every 1 percent increase in unemployment gap (that is, when actual unemployment exceeds its natural rate by 1 percent), the output gap is -2%, (that is, actual real GDP is less than its potential level by 2 percent).

2015: (7840-8000)*100/8000 = -2*(u - 6) or -2 /-2 = u - 6. This gives unemployment rate = 7%.

2016: unemployment gap = 0 and so potential GDP = actual GDP = 8100

2017: unemployment gap = 0.5% so we have (Y - 8200)*100/8200 = -2*0.5 gives actual GDP = 8118

2017:GDP gap = (8415-8250)*100/8250 = 2% and so 2 = -2(5 - u*) which gives natural rate of unemployment = 6%.